Please answer both questions to the best of your ability! It would be GREATLY APPRECIATED.
10: Water has a specific heat of 4.18 1/g C. If 73.7 g of water at 87.9C loses 2.6 k of heat, what is the final temper Q11: A 59.7 g piece of metal that had been submerged in boiling water was quickly transferred into 0.060 L of water initially at 22.0 *C The final temperature is 28.5 .?. Use these data to determine the specific heat of the metal. A specific information about water is available in the previous question.
Answer #1

Answer 10 :-

Let the final temperature be =TºC

Then,

?T=(87.9?T)ºC

Mass of water, m=0.0737kg

Heat lost, H=2.6kJ

specific heat of water, s=4.18kJkg?1ºC?1

Therefore,

H=ms?T

2.6 = 0.0737 * 4.186 * (87.9?T)

87.9 - T = 2.6 / (0.0737 * 4.186)

87.9 - T = 8.43

T = 87.9 - 8.43

T = 79.5 °C